OpenStudy (gokuporter):

Point C is graphed on a number line at –7. Point D is 13 units away from point C on the number line. What could be the coordinates of point D? Choose all answers that are correct. A. –20 B. 3 C. 6 D. 20

OpenStudy (anonymous):

two numbers can be point C -7 -13 = and -7 + 13 =